# Fine-Grained Motion Compression and Selective Temporal Fusion for Neural B-Frame Video Coding

### v1

@@ IV-A2 Evaluation Datasets
The compression performance is evaluated on three standard benchmark datasets: the MCL-JCV dataset [58] containing 30 diverse 19201080 sequences, the UVG dataset [59] with 7 challenging 1080p large-motion sequences, and the HEVC common test sequences [60] comprising 16 videos (Classes B, C, D, E) at resolutions spanning from 416240 to 19201080. We compress 96 (2 complete GOPs and 1 incomplete GOP) and 97 (3 complete GOPs) consecutive frames of each test sequence to verify the generalization of our codec for complete GOP and incomplete GOP. The detailed GOP structures can be referred to [33].
@@ IV-A3 Implementation Details
During training, the Lagrange multiplier in the loss function is configured as (85, 170, 380, 840) to measure the relative importance of distortion and bit rate. At inference time, our codec achieves different rate points by adjusting its learned quantization steps. The implementation is built in PyTorch with CUDA acceleration, trained using the AdamW optimizer [61] with default hyperparameter settings. The batch size is set to 8.
@@ IV-A4 Evaluation Metrics
We compute Peak Signal-to-Noise Ratio (PSNR) for quality assessment, and use bits per pixel (bpp) to measure the bitrate cost. For comparative analysis between different codecs, we utilize the Bjøntegaard-Delta rate (BD-rate) metric, which calculates the average bitrate difference (%) relative to an anchor codec. Negative BD-rate values indicate bitrate savings, positive values denote increased bitrate requirements.
@@
IV-B Quantitative Evaluation Results
IV-B1 Comparison with Existing Neural Video Codecs
@@ IV-B2 Comparison with Traditional Video Codecs
We further compare with traditional video codecs. Specifically, we include HM-16.20 [62] (H.265/HEVC reference software) tested under random-access (encoder_randomaccess_main_rext, HM-RA-GOP16) and low-delay (encoder_lowdelay_vtm, HM-LDB-GOP4) configurations, along with VTM-13.2 (H.266/VVC reference software) evaluated under random-access (encoder_randomaccess_vtm_gop16, VTM-RA-GOP16 and encoder_randomaccess_vtm, VTM-RA-GOP32) and low-delay (encoder_lowdelay_vtm, VTM-LDB-GOP8) configurations. All codecs are configured with an intra-period of 32. When compressing 96 frames, as reported in Fig. 6 and Table I, our codec achieves comparable performance to VTM, with an average BD-rate reduction of –34.5% versus –34.1% (VTM-RA-GOP16) and –35.8% (VTM-RA-GOP32). When testing 97 frames, our codec even achieves higher performance than VTM, with an average BD-rate reduction of –38.0% versus –32.7% (VTM-RA-GOP16) and –35.7% (VTM-RA-GOP32). These results demonstrate the strong potential of neural B-frame codecs.
@@ IV-D Runtime and Computational Complexity Comparison
| Model | IDMAE | PBAQS | IMEM | BWF | HIA | BD-rate |
| ✗ | ✗ | ✗ | ✗ | ✗ | 0.0 | |
| ✓ | ✗ | ✗ | ✗ | ✗ | –2.6 | |
| ✓ | ✓ | ✗ | ✗ | ✗ | –4.9 | |
| ✓ | ✓ | ✓ | ✗ | ✗ | –6.3 | |
| ✓ | ✓ | ✓ | ✓ | ✗ | –9.1 | |
| ✓ | ✓ | ✓ | ✓ | ✓ | –10.4 |
@@ IV-E1 Effectiveness of Fine-Grained Motion Compression Method
We first conduct an ablation study to validate the effectiveness of our proposed fine-grained motion compression method. To reduce training time, we only use 7-frame video sequences to train the models in the ablation studies, instead of fine-tuning them with the long-sequence dataset [33]. Our ablation study shows substantial performance gains through progressive integration of key components, as evidenced in Table IV. The interactive dual-branch motion auto-encoder (IDMAE) forms the foundation with a 2.6% BD-rate improvement, while the per-branch adaptive quantization steps (PBAQS) further deliver a 2.3% gain (from –2.6% to –4.9% BD-rate) by enabling fine-grained rate allocation tailored to each motion direction. The interactive motion entropy model (IMEM) further brings another 1.7% gain (from –4.9% to –6.3% BD-rate), which effectively captures the inherent correlations between bi-directional motion latent representations for enhanced compression efficiency. We compare the motion coding costs of the model with and without our proposed method in Fig. 9. We can observe that our method can effectively reduce the bi-directional motion compression bitrate.
@@ IV-E2 Effectiveness of Selective Temporal Fusion Method
Similarly, we continue to conduct another ablation study to validate the effectiveness of our selective temporal fusion method by progressively integrating its key components into the baseline Model . As presented in Table IV, the introduction of bi-directional weighting fusion (BWF) first improves the compression performance by 2.8% BD-rate reduction (from –6.3% to –9.1% BD-rate), followed by an additional 1.3% gain (reaching –10.4% BD-rate) when incorporating the hyperprior-based implicit alignment (HIA).
@@
V Conclusion
We first propose a fine-grained motion compression method for neural B-frame video coding. By leveraging an interactive dual-branch motion auto-encoder, per-branch adaptive quantization steps, and an interactive motion entropy model, the bi-directional motion coding costs are effectively reduced. Furthermore, we propose a selective temporal fusion method. Using bi-directional fusion weights and hyperprior-based implicit alignment, our method enables discriminative utilization of bi-directional temporal contexts and temporal priors in the contextual encoder-decoder and contextual entropy model. Experimental results verify that our proposed codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configuration. In the future, we target to achieve a better rate-distortion-complexity trade-off for neural B-frame coding.

### v2

@@ IV-A2 Evaluation Datasets
The compression performance is evaluated on three standard benchmark datasets: the MCL-JCV dataset [64] containing 30 diverse 19201080 sequences, the UVG dataset [42] with 7 challenging 1080p large-motion sequences, and the HEVC common test sequences [3] comprising 16 videos (Classes B, C, D, E) at resolutions spanning from 416240 to 19201080. We compress 96 (two complete GOPs and one incomplete GOP) and 97 (three complete GOPs) consecutive frames of each test sequence to verify the generalization of our codec for complete GOP and incomplete GOP. The detailed GOP structures can be referred to [57].
@@ IV-A3 Implementation Details
During training, the Lagrange multiplier in the loss function is configured as (85, 170, 380, 840) to measure the relative importance of distortion and bit rate. At inference time, our codec achieves different rate points by adjusting its learned quantization steps. The implementation is built in PyTorch with CUDA acceleration, trained using the AdamW optimizer [24] with default hyperparameter settings. The batch size is set to 8. The learning rate is gradually reduced from an initial value of to a final value of throughout the training process.
@@ IV-A4 Evaluation Metrics
Following previous works [26, 27, 28, 29, 57, 44, 17, 7, 56, 54], we compute Peak Signal-to-Noise Ratio (PSNR) in RGB color space for quality assessment, and use bits per pixel (bpp) to measure the bitrate cost. For comparative analysis between different codecs, we utilize the Bjøntegaard-Delta rate (BD-rate) metric, which calculates the average bitrate difference (%) relative to an anchor codec. Negative BD-rate values indicate bitrate savings, positive values denote increased bitrate requirements.
IV-B Quantitative Evaluation Results
IV-B1 Comparison with Existing Neural Video Codecs
It is worth noting that while our method establishes a new state-of-the-art in terms of compression performance, we recognize the distinct advantages of other innovative approaches. For instance, the motion-free concept explored in [43] and the fast online motion resolution adaptation in [44] present compelling alternative pathways. These ideas, focusing on complexity reduction and operational flexibility, offer valuable insights for future work aimed at achieving a better trade-off between performance, complexity, and practicality in neural video coding.
@@ IV-B2 Comparison with Traditional Video Codecs
We further compare with traditional video codecs. Specifically, we include HM-16.20 [16] (H.265/HEVC reference software) tested under random-access (encoder_randomaccess_main_rext, HM-RA-GOP16) and low-delay (encoder_lowdelay_vtm, HM-LDB-GOP4) configurations, along with VTM-13.2 (H.266/VVC reference software) evaluated under random-access (encoder_randomaccess_vtm_gop16, VTM-RA-GOP16 and encoder_randomaccess_vtm, VTM-RA-GOP32) and low-delay (encoder_lowdelay_vtm, VTM-LDB-GOP8) configurations. As the PSNR is calculated in RGB colorspace, we follow the testing pipelines of [26, 27, 28, 29, 57, 44, 17, 7, 56, 54], and convert RGB video frames to YUV444 sequences. For better compression performance of traditional codecs, we set their internal colorspace as YUV444. All codecs are configured with an intra-period of 32. When compressing 96 frames, as reported in Fig. 6 and Table I, our codec achieves comparable performance to VTM, with an average BD-rate reduction of –34.5% versus –34.1% (VTM-RA-GOP16) and –35.8% (VTM-RA-GOP32). When testing 97 frames, our codec even achieves higher performance than VTM, with an average BD-rate reduction of –38.0% versus –32.7% (VTM-RA-GOP16) and –35.7% (VTM-RA-GOP32). These results demonstrate the strong potential of neural B-frame codecs.
It is important to acknowledge that while our results highlight the potential of neural video coding, the field stands on the shoulders of decades of progress in traditional video coding. The sophisticated tools and profound insights developed for standards like H.264/AVC, H.265/HEVC, and H.266/VVC, such as advanced motion estimation methods [22], represent a vast repository of knowledge. These well-established principles continue to offer invaluable inspiration and serve as a crucial benchmark for guiding the future development of more efficient, robust, and practical neural video codecs.
@@ IV-D Runtime and Computational Complexity Comparison
| Model | IDMAE | PBAQS | IMEM | BWF | HIA | BD-rate | MACs/pixel |
| ✗ | ✗ | ✗ | ✗ | ✗ | 0.0 | 0.0 | |
| ✓ | ✗ | ✗ | ✗ | ✗ | –2.6 | 311.93K | |
| ✓ | ✓ | ✗ | ✗ | ✗ | –4.9 | 311.93K | |
| ✓ | ✓ | ✓ | ✗ | ✗ | –6.3 | 331.88K | |
| ✓ | ✓ | ✓ | ✓ | ✗ | –9.1 | 357.80K | |
| ✓ | ✓ | ✓ | ✓ | ✓ | –10.4 | 377.73K |
@@ IV-E1 Effectiveness of Fine-Grained Motion Compression Method
We first conduct an ablation study to validate the effectiveness of our proposed fine-grained motion compression method. To reduce training time, we only use 7-frame video sequences to train the models in the ablation studies, instead of fine-tuning them with the long-sequence dataset [57]. Our ablation study shows substantial performance gains through progressively integrating the key components into the baseline model , as evidenced in Table IV. The interactive dual-branch motion auto-encoder (IDMAE, ) forms the foundation with a 2.6% BD-rate improvement, while the per-branch adaptive quantization steps (PBAQS, ) further deliver a 2.3% gain (from –2.6% to –4.9% BD-rate) by enabling fine-grained rate allocation tailored to each motion direction. The interactive motion entropy model (IMEM, ) further brings another 1.4% gain (from –4.9% to –6.3% BD-rate), which effectively captures the inherent correlations between bi-directional motion latent representations for enhanced compression efficiency. We compare the motion coding costs of the model with and without our proposed method in Fig. 9. We can observe that our method can effectively reduce the bi-directional motion compression bitrate.
Regarding the computational complexities of these methods, it is important to note that IDMAE, PBAQS, and IMEM are inherently coupled components of our fine-grained motion compression method. The IDMAE method () establishes the fundamental dual-branch structure necessary for processing bi-directional motions separately, which naturally introduces additional 311.93K MACs/pixel computational complexity increase compared to the single-branch baseline (). Without this foundational structure, the subsequent PBAQS and IMEM components would not be feasible. The PBAQS method () maintains the same computational complexity as but increases the model parameters slightly by employing independent quantization steps for each branch instead of the parameter-shared steps used in . The IMEM method () adds 19.95K MACs/pixel by incorporating cross-directional dependencies in the motion entropy modeling.
@@ IV-E2 Effectiveness of Selective Temporal Fusion Method
Similarly, we continue to conduct another ablation study to validate the effectiveness of our selective temporal fusion method by progressively integrating its key components into the model . As presented in Table IV, the introduction of bi-directional weighting fusion (BWF, ) first improves the compression performance by 2.8% BD-rate reduction (from –6.3% to –9.1% BD-rate), followed by an additional 1.3% gain (reaching –10.4% BD-rate) when incorporating the hyperprior-based implicit alignment (HIA, ). In terms of computational complexity, the BWF method () introduces additional 25.92K MACs/pixel computational complexity compared to , while the HIA method () further adds 19.93K MACs/pixel. The relatively modest increase in complexity shows that our proposed selective temporal fusion method achieves effective performance improvement with limited computational overhead.
@@
V Conclusion and Future Work
We first propose a fine-grained motion compression method for neural B-frame video coding. By leveraging an interactive dual-branch motion auto-encoder, per-branch adaptive quantization steps, and an interactive motion entropy model, the bi-directional motion coding costs are effectively reduced. Furthermore, we propose a selective temporal fusion method. Using bi-directional fusion weights and hyperprior-based implicit alignment, our method enables discriminative utilization of bi-directional temporal contexts and temporal priors in the contextual encoder-decoder and contextual entropy model. Experimental results verify that our proposed codec outperforms state-of-the-art neural B-frame codecs and achieves comparable or even superior compression performance to the H.266/VVC reference software under random-access configuration.
While the proposed codec achieves impressive compression performance, its increased computational complexity relative to existing neural codecs presents a key challenge for practical deployment. Our future work will, therefore, focus on achieving a more favorable complexity-performance trade-off. A promising direction, inspired by motion-free neural B-frame coding approaches, is the development of a hybrid coding framework. Specifically, within a hierarchical B-frame structure, frames at higher temporal layers exhibit smaller motion due to their proximity to the references. For these frames, we plan to leverage lightweight, interpolation-based prediction to generate temporal contexts, effectively bypassing the computationally intensive processes of motion estimation, compression, and compensation. We anticipate that this content-adaptive strategy will significantly reduce the average computational cost while retaining the majority of the compression performance gains achieved by our current design.